1
(1 point) Archeologists have discovered a sample of ancient Egyptian skulls from 4000 B.C. A sample of 6 Egyptian skull bread 2
Answer #1

Solution(a)
Sample mean can be calculated as
Sample mean (Xbar) = \sum Xi/n = (135+127+138+127+128+141+130+128+132+134)/10 = 1320/10 = 132
Sample standard deviation can be calculated as
Sample standard deviation S = sqrt(\sum(Xi-mean)^2/(n-1)) = sqrt((135-132)^2 + (127-132)^2 + (138-132)^2 + (127-132)^2 + (128-132)^2 + (141-132)^2 + (130-132)^2 + (128-132)^2 + (132-132)^2 + (134-132)^2)/9)

X (Xi-mean) (Xi-mean)^2
135 3 9
127 -5 25
138 6 36
127 -5 25
128 -4 16
141 9 81
130 -2 4
128 -4 16
132 0 0
134 2 4


Sample standard deviation (S) = sqrt((9+25+36+25+16+81+4+16+0+4)/9) = sqrt(216/9) = 4.8990
Solution(b)

Sample mean can be calculated as
Sample mean (Xbar) = \sum Xi/n = (136+124+137+122+120+119)/6 = 758/6 = 126.33
Sample standard deviation can be calculated as
Sample standard deviation S = sqrt(\sum(Xi-mean)^2/(n-1)) = sqrt((136-126.33)^2 + (124-126.33)^2 + (137-126.33)^2 + (122-126.33)^2 + (120-126.33)^2 + (119-126.33)^2)/5) = sqrt(325.3334/5) = 8.0664

X (Xi-mean) (Xi-mean)^2
136 9.67 93.5089
124 -2.33 5.4289
137 10.67 113.8489
122 -4.33 18.7489
120 -6.33 40.0689
119 -7.33 53.7289